SUMMER 2009 UNIVERSITY OF CHICAGO REU PROGRAM Apprentice 5 week program

CHECK ONE:

I will participate in the program; I understand that it is my responsibility to be present M-F 9:00-5:00 from June 22 to July 17, 2009 and that I will not engage in any other employment during that period without the explicit approval of the program director. The program will be fully operational through July 24, and attendance for five weeks is very strongly recommended although not required. I also understand that I am welcome to attend the program through August 14 and that a final short piece of mathematical writing (expository or research) is required.

I will not participate in the program

NAME:

ADDRESS:

E-MAIL: